5. Laplace applications (1) 5*e**tcostdt = (2) [sin 21 dt (3) [** :[e* sin Ardt Can you show steps.
Answer #1

P Jetcost at {{ë FLt)} = f(s+a) = f(542) L{F(t)} = {{t6ott} - - {f(s)} #18)= 4{cost} ) = -f($*+1)x1 = 5(23)? (527192 (s+2) ²-내쁜 Life (\) 3 - Dayl ) . | - Lf Eter} = (+3) | | L{ sinst3 = () . . fancer = . fronurs = t ) 1305) - it | F(St 3) = Es316 ©9)